E6-4 Analyzing Changes in Price, Cost Structure, Degree of Operating Leverage [LO 6-4, 6-5] Cove’s Cakes is a local bakery. Price and cost information follows: Price per cake $ 14.11 Variable cost per cake Ingredients 2.29 Direct labor 1.16 Overhead (box, etc.) 0.11 Fixed cost per month $ 3,376.00 Required:

1. Calculate Cove’s new break-even point under each of the following independent scenarios: (Round your answer to the nearest whole number.)

a. Sales price increases by $2.00 per cake.

b. Fixed costs increase by $475 per month.

c. Variable costs decrease by $0.29 per cake.

d. Sales price decreases by $0.60 per cake.

2. Assume that Cove sold 330 cakes last month. Calculate the company’s degree of operating leverage. (Do not round intermediate calculations. Round your answer to 2 decimal places.)

3. Using the degree of operating leverage calculated in Requirement 2, calculate the change in profit caused by a 14 percent increase in sales revenue. (Round your final answer to 2 decimal places (i.e. .1234 should be entered as 12.34%.))
Answer #1

Current sales price per cake: $ 14.11

Current variable expenses per cake = $ 3.56

Current contribution margin per cake = $ ( 14.11 - 3.56 ) = $ 10.55

Current break-even point in units = $ 3,376 / $ 10.55 = 320 units.

1. a. Sales price increase by $ 2 per unit :

Contribution margin per unit = $ 12.55

Break-even point = $ 3,376 / $ 12.55 = 269 units

Break-even sales dollars = 269 units x $ 16.11 = $ 4,333.59

b. Fixed cost increases by $ 475 per month :

Break -even point in units = $ ( 3,376 + 475 ) / $ 10.55 = 365.02 units

Break-even point in sales dollars = 365.02 x $ 14.11 = $ 5,150.43

c. Variable cost decrease by $ 0.29 per cake.

New contribution margin = $ 10.55 + $ 0.29 = $ 10.84

New break-even point = $ 3,376 / $ 10.84 = 311.44 units

New break-even point in sales dollars = 311.44 units x $ 14.11 = $ 4,394.42

d. Sales price decrease by $ 0.60 per cake:

New contribution margin = $ 10.55 - $ 0.60 = $ 9.95

Break-even point in units = $ 3,376 / $ 9.95 = 339.30 units.

Break-even sales dollars = 339.30 x $ (14.11 - 0.60) = $ 4,583.94

2. Degree of Operating Leverage : 33

Sales Revenue ( 330 x $ 14.11 ) $ 4,656.30
Variable Expenses ( 330 x $ 3.56 ) 1.174.80
Contribution Margin 3,481.50
Fixed Expenses 3,376
Net Operating Income ( loss ) $ 105.50

Degree of Operating Leverage = Contribution Margin / Net Operating Income = $ 3,481.50 / $ 105.50 = 33

3. Change in profit = 14 % x 33 = 462 %